The following Information is avallable for Quality Book Sales sales on account and accounts recelvable: Accounts receivable balance, January 1, Year 2 Allowance for doubtful accounts, January 1, Year 2 Sales on account, Year 2 Collection on accounts receivable, Year 2 $ 78,800 4,778 553,808 559,8e0 After several collection attempts, Quality Book Sales wrote off $2720 of accounts that could not be collected. Quality Book Sales estimates that 5 percent of the ending accounts receivable balance will be uncollectible. Required a. Compute the following amounts: (1) Using the allowance method, the amount of uncollectible accounts expense for Year 2 (2) Net realizable value of recelvables at the end of Year 2. Amount of uncollectible accounts Net realizable valueb. Record the general journal entries to (1) Record sales on account for Year 2 (2) Record cash collections from accounts recelvable for Year 2. (3) Write off the accounts that are not collectible 4) Record the estimated uncollectible accounts expense for Year 2 (If no entry Is required for a transaction/event, select No Journal entry required In the first account field.) View transaction list Journal entry worksheet Record sales on account for Year 2. Note: Enter debits before credits, Event General Journal Debit Credit Record entry Clear entry View general jourmal
Answer #1

1a) Uncollectible account expenses = 553000*5% = $27650

1b) Calculate balance :

Allowance for doubtful accounts = (4770+27650-2720) = $29700

Net realizable value = (78800+553000-559000-2720)-29700 = $40380

2) Journal entries

No account and explanation debit credit
1 Account receivable 553000
Sales revenue 553000
(To record sales)
2 Cash 559000
Account receivable 559000
(To record collection)
3 Allowance for doubtful accounts 2720
Account receivable 2720
(To record write off)
4 Bad debt expense (553000*5%) 27650
Allowance for doubtful accounts 27650
(To record uncollectible account expense)